OpenStudy (raiders88):

solve the system of equations by substitutions : x=-4y x+5y

OpenStudy (raiders88):

@misty1212

OpenStudy (niahloves):

10y

OpenStudy (lynfran):

u r given x=-4y where ever u see x sub that into x+5y

OpenStudy (lynfran):

what did u get?

OpenStudy (raiders88):

idk

OpenStudy (lynfran):

-4y+5y

OpenStudy (lynfran):

simplify

OpenStudy (lynfran):

is that even the equations?

OpenStudy (lynfran):

x=-4y x+5y ??

OpenStudy (niahloves):

9y?

OpenStudy (lynfran):

-4y+5y equals y its like saying 5y-4y....

OpenStudy (raiders88):

plz just tell me what it is

OpenStudy (raiders88):

@misty1212

OpenStudy (raiders88):

@welshfella

OpenStudy (welshfella):

the second part is not an equation ( no '=')

OpenStudy (welshfella):

x + 5y is not an equation

OpenStudy (lynfran):

thats what i realize..and why i asked "is that even the equations"

OpenStudy (mathstudent55):

A system of equation needs to have at least two equations. x = -4y is an equation. Where is the second equation?

OpenStudy (dinamix):

-4y+5y = 5y-4y= y(5-4) = y @raiders88